### 08 Jan 2024 v1

@@ 2.1 Problem formulation and high fidelity discretization
Given the grain microstructure, we compute several quantities of interest (QoIs): number (or percentage) of eliminated grains; grain size distribution; and volume-averaged misorientation. The number of eliminated grains is relevant in epitaxial growth [54]. We compute $n_{G}=n_{G}(z)$, the accumulated grain eliminations from the initial interface location to a height $z$ (see Fig. 1f). Grain size distribution is a common descriptor of grain shape statistics and it affects mechanical properties such as strength and ductility. Grain size $d$ of grain $g$ is defined by its volume-equivalent diameter $d_{g}=(6\mathcal{V}_{g}/\pi)^{1/3}$, where $\mathcal{V}_{g}$ is the grain volume. Fig. 1d shows the probability distributions of the grain size for Fig. 1a and Fig. 1b. Volume-weighted misorientation $\Delta\boldsymbol{\theta}$ quantifies the alignment of the polycrystalline orientation with the prescribed temperature gradient direction $z$. $\Delta\boldsymbol{\theta}=\sum_{g}\mathcal{V}_{g}\boldsymbol{\theta}_{z,g}/\sum_{g}\mathcal{V}_{g}$ and $\boldsymbol{\theta}_{z,g}$ is the angle between the crystal orientation of grain $i$ and the $z$-axis. As shown in Fig. 1f, $\Delta\boldsymbol{\theta}$ decreases as the more aligned grains out-compete the misaligned grains. The discretization of Eq. 2.1 is discussed in Section 2.8. inline]Explain that we need several simulations due to random initial conditions in order to obtain expectations of the observed variables
@@
Fig. 1 shows an example of a simulation with domain size $(120\mu m,120\mu m,50\mu m)$. The interface width $W_{0}=0.1\mu m$ and the mesh size $dx=0.08\mu m$ [6]. Therefore the grid size is $1500\times 1500\times 625$. We use no-flux boundary conditions at the top and bottom surface ($z=0,z=L^{\prime}_{z}$) and periodic boundary conditions at the four sides of the domain. The initial conditions for grain microstructure in the substrate are generated by generating random points and using a Voronoi diagram with periodic boundary conditions. In Fig. 1a, the substrate has 900 grains. The orientation of each grain is sampled from the unit sphere. If $\boldsymbol{d}$ is the vector representing the grain’s orientation, we first sample $\boldsymbol{d}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and set $\boldsymbol{d}=\boldsymbol{d}/|\boldsymbol{d}|_{2}$. In this simulation $G_{z}$ and $R_{z}$ are $10K/\mu m$ and $2m/s$, respectively. Fig. 1b is the resulting phase field microstructure after 10K time steps, which required approximately one hour on one Nvidia A100 GPU.
@@ 2.2 GrainGNN
The input to GrainGNN is the initial state $\boldsymbol{\xi}$ of the substrate (see Fig. 1a), process parameters $\boldsymbol{p}=\{G_{z},R_{z}\}$, and a domain size $L=(L_{x},L_{y},L_{z})$. $\boldsymbol{\xi}$ is parameterized by the coordinates of the Voronoi vertices and orientations of grains: $\boldsymbol{\xi}=\{x_{j},y_{j}\}^{n_{j}}_{j=1}\cup\{\boldsymbol{\theta}_{g}\}^{n_{g}}_{g=1}$, where $n_{j}$ is the number of Voronoi vertices. The output of GrainGNN is the grain orientation field $\boldsymbol{\theta}(x,y,z)$ in a domain with size $L$. Notice that here $L_{z}$ is the height of the final polycrystal (as shown in Fig. 1b); for phase field simulations we let $L^{\prime}_{z}>L_{z}$ to avoid the effect of the top boundary. In sections Section 2.3-Section 2.6 we explain the basic ingredients of GrainGNN.
  • Image-to-graph step. We start with a spatial field compression method to reduce the large dimensionality of the phase field data. We define a planar graph $G$ and hand-crafted features $F$ to represent a layer of grain microstructure (see Fig. 2). Do define these features from $\{\phi_{\alpha}\}_{\alpha=1}^{n_{g}}$ we first combine the fields to to a single image $\boldsymbol{\theta}(x,y,z)$ the grain orientation at each point $x,y,z$. Then we propose an image-to-graph algorithm $\phi_{\alpha}(x,y,z)\rightarrow(G_{l},F_{l}),\ l=0,1,2,...$ to extract vertices and edges of grains at cross-sections defined at $z_{l}$. The features contain geometric information about the size of the grains and their boundary. The definitions of the image, graph, and features are discussed in Section 2.3.
  • Graph evolution step. We now train a network for the map $(G_{l-1},F_{l-1})\rightarrow(G_{l},F_{l})$. The challenge is that graphs $G_{l-1}$ and $G_{l}$ can be topologically different due to grain or/and edge elimination events. To address this we need to predict both vertex feature evolution $\Delta F$ as well as a set of grain neighbor switching events $\mathcal{S}_{E}$ and a set of grain elimination events $\mathcal{S}_{G}$. We discuss the graph evolution in detail in Section 2.4. The prediction $(G_{l},F_{l})$ is done by first using GrainGNN to predict $\Delta F$ and probabilities of events and then using a an algorithm reconstruct $G_{l}$ and its features.
@@ 2.3 Graph representation of microstructure
Now, let us describe the compressed representation of the microstructure using hand-crafted features. Consider a solution of Eq. 2.1. From this simulation we sample $n_{l}$ layers of field data as training data. We define a “layer” at height $z_{l}$ as the microstructure at the time $t_{l}$ when the solid-liquid interface (SLI) reaches $z_{l}$ (see Fig. 2a). More precisely, $t_{l}$ is defined as the time at which the lowest point (order in $z$ coordinate) at the SLI reaches $z_{l}$.
@@
We solve Eq. 2.1 to generate training data and to verify the predictions of the surrogate. Although a simulation can have thousands of time steps, we subsample to $n_{l}$ planes using equispaced sampling for $z_{l}$ with height increment $\Delta z$, so $z_{l}=l\Delta z$. Typically we use $n_{l}=20$. We discuss the choice $\Delta z$ and $n_{l}$ in Section 2.4.
@@
Once $I_{l}$ have been identified, who extract graph $(G_{l},F_{l}$ (Fig. 2c) as follows. $G(V,E)$ is an undirected graph where $V$ and $E$ denote the sets of vertices and edges respectively. In GrainGNN, $G$ has two types of vertices $V=V_{g}\cup V_{j}$. $V_{g}=\{{g}_{1},g_{2},...,{g_{n_{g}}}\}$ represents grains. $V_{\boldsymbol{j}}=\{{j}_{1},{j}_{2},...,{j_{n_{j}}}\}$ are junction vertices, where $n_{j}$ is the number of junctions. A junction vertex refers to a point in image $I_{l}$ that has three grain neighbors (e.g., $j_{1}$ in Fig. 2d). When we create the initial condition of the substrate with a Voronoi diagram, the Voronoi vertices are also the junction vertices of $G$. $G$ has two types of undirected edges $E=E_{jj}\cup E_{jg}$. $E_{jj}$ are the edges between junction vertices and can be thought of as a representation of actual inter-grain boundaries on the $x$-$y$ plane. $E_{jg}$ are the junction-grain edges (dashed lines in Fig. 2e).
@@
As we will see later to define we need to define a maximum degree (number of edges) for a junction vertex. The maximum number of edges for a junction vertex depends on the physical setting in the problem. In rapid solidification of metals, the grain boundary anisotropy is relatively weak compared to kinetic anisotropy [55], and a triple junction is more stable than a quadruple junction [19]. Thus we only consider triple junctions in this work, which means a junction vertex can only connect to the other three junction vertices and three grain vertices. In this case, $n_{j}=2n_{g}$, $|E_{jj}|=3n_{g}$, and $|E_{jg}|=6n_{g}$. Our handling of the graph topological changes discussed in Section 2.4 is based on this simplification of the graph structure.
@@
Let’s now discuss how we define the vertices and edges from image $I_{l}$. We now present the method to extract $G_{l}$ and $F_{l}$ from $I_{l}$. First, we identify the junction locations $(x_{j}$, $y_{j})$. For each pixel $(x,y)$ of $I_{l}$, we inspect its eight neighbors each associated with a grain index. If there are three distinctive grain indices in eight neighbors, we consider pixel $(x,y)$ to be a junction pixel, for example, $j_{1}$ in Fig. 2d. Every junction $j$ is associated with a unique triplet of grain indices $N_{j}=\{g_{1},g_{2},g_{3}\}$. Commonly there are multiple adjacent pixels that have the same triplet; we keep the one pixel whose eight neighbors are more evenly distributed to the three grain indices. Then using the triplet indices $N_{j}$ we add three edges to the $E_{jg}$ set: e.g., for $j_{1}$, the $E_{jg}$ edges are $({j_{1}},g_{1})$, $({j_{1}},g_{2})$, $({j_{1}},g_{3})$. We create $E_{jj}$ edges by $E_{jj}=\{(j_{1},j_{2}):N_{j_{1}}\cap N_{j_{2}}=2,\forall j_{1},j_{2}\in G\}$. For example in Fig. 2e, edge $(j_{1},j_{2})$ exists because $j_{1}$ and $j_{2}$ share the same neighboring grains $g_{1}$ and $g_{2}$. For $x_{g}$, the coordinates of grain vertex $g$, we average the coordinates of its junction neighbors $x_{g}=\sum x_{j,k}/|N_{g}|,\ \forall k\in N_{g}$. $N_{g}$ is the set of junctions connected to grain $g$.
@@
We clarify that these are input features to the LSTM and serve as a reduced representation of the phase field. They are not the network hidden features, which are of course determined during training. We define vertex and edge features as follows:
@@
    where $x_{j}$, $y_{j}$, and $z_{l}$ are the 3D coordinates of a junction vertex; $\Delta x_{j}$ and $\Delta y_{j}$ are in-plane displacements of junctions with respect to their locations in the previous layer $G_{l-1}$, i.e., $\Delta x_{j}=x_{j,l}-x_{j,l-1},\Delta y_{j}=y_{j,l}-y_{j,l-1}$. $\Delta z=z_{l}-z_{l-1}$ is the distance between two sampled images. For $l=0$, we set $\Delta x_{j}=\Delta y_{j}=\Delta z=0$. $G_{z}$ and $R_{z}$ here are constants; they’re repeated at each vertex and allow them to vary for each junction location.
@@
    where $x_{g}$, $y_{g}$, and $z_{l}$ are the coordinates of the grain vertices; $s_{g}$ is the grain cross-sectional area; and $v_{g}$ is the grain excess volume above the interface height $z_{l}$ [17]. $\boldsymbol{\theta}_{z}$ is the angle between the $z$-axis and the preferred growth direction ¡100¿; $\boldsymbol{\theta}_{x}$ is the angle between the $x$-axis and the projection of ¡100¿ direction on the $x$-$y$ plane. $\Delta s_{g}=s_{g,l}-s_{g,l-1}$ is the change of cross-sectional area; $\Delta s_{g}=0$ at $l=0$. In Eq. 2.5 and Eq. 2.4, $\boldsymbol{f}^{0}_{j}$ and $\boldsymbol{f}^{0}_{g}$ are constant vectors to normalize features to [0, 1]. $\boldsymbol{f}^{0}_{j}=\begin{bmatrix}L_{x},L_{y},L_{z},G_{\mathrm{max}},R_{\mathrm{max}},L_{x},L_{y},L_{z}\end{bmatrix}$ and $\boldsymbol{f}^{0}_{g}=\begin{bmatrix}L_{x},L_{y},L_{z},L_{x}L_{y},L_{x}L_{y}L_{z},1,1,1,1,L_{x}L_{y},L_{z}\end{bmatrix}$. $G_{\mathrm{max}}$ and $R_{\mathrm{max}}$ are maximum values of $G_{z}$ and $R_{z}$ in the training data.
@@
Fig. 3: Topological events with the update of graph connectivity. (a) A neighbor-switching event requires changing the graph topology: an $O_{E}$ operation. Grains $g_{3}$ and $g_{4}$ initially are neighbors. After applying $O_{E}$, $g_{1}$ and $g_{2}$ are neighbors. From \@slowromancapi@ to \@slowromancapii@, $j_{1},j_{2}$ merge to a quadruple junction; from \@slowromancapii@ to \@slowromancapiii@, the quadruple junction splits into two triple junctions. This event replaces edges $(j_{1},j_{5})$, $(j_{2},j_{4})$ (in red color) with the edges $(j^{\prime}_{2},j_{5})$, $(j^{\prime}_{1},j_{4})$ (in blue color). The total number of edges remains the same. (b) A grain elimination event. From \@slowromancapi@ to \@slowromancapii@, a neighbor-switching event occurs on $(j_{1},j_{2})$; $g_{3}$ become a 2-side grain. \@slowromancapii@ to \@slowromancapiii@ is an $O_{G}$ operation. It removes $g_{3}$, $j^{\prime}_{1}$, $j_{3}$, a couple of edges, and replace them with $(j^{\prime}_{2},j_{4})$.
@@ 2.4 Graph evolution
Grain neighbor switching events: Grain neighbor switching happens because a grain edge gets smaller and eventually disappears. A disappearing edge creates a four-edge junction (Fig. 3a). Since in our PDE model four-edge junctions are unstable, a new edge starts growing and the four-junction becomes a three junction. This new edge causes the neighbor switch. Physically, a neighbor-switching event involves four grains, where two grains lose one face and the other two grains gain one face. From \@slowromancapi@ to \@slowromancapii@, grain $g_{1}$ and $g_{2}$ push the vertices $j_{1}$ and $j_{2}$ to move toward each other. At some intermediate time, $j_{1}$ and $j_{2}$ merge to form a quadruple junction. Then from \@slowromancapii@ to \@slowromancapiii@, the quadruple junction separates into two new junctions $j^{\prime}_{1}$ and $j^{\prime}_{2}$, which replace $j_{1}$ and $j_{2}$. This event changes several edges, grain $g_{1}$ and $g_{2}$ grain an additional junction-grain edge; $g_{3}$ and $g_{4}$ lose one junction-grain edge; $(j^{\prime}_{2},j_{5})$ and $(j^{\prime}_{1},j_{4})$ replace $(j_{1},j_{5})$ and $(j_{2},j_{4})$; the total number of edges remain the same. The grain neighbor-switching event is characterized by the removal of $(j_{1},j_{2})$.
@@
Matching $G_{l-1}$ and $G_{l}$ and handling elimination events. Our graph update requires computing $\Delta F,\mathcal{S}_{E},\mathcal{S}_{G}$. We define
@@
Neighbor-switching events are harder to handle: they result in junctions whose triplets $N_{j}$ are different in $G_{l}$ and $G_{l-1}$. Let $\mathcal{U}_{l-1}$ be the junction vertices in $G_{l-1}$ but not matched in $G_{l}$, $\mathcal{U}_{l-1}=\{j:j\in G_{l-1}\wedge j\notin G_{l}\}$. We enumerate every two junctions in $\mathcal{U}_{l-1}$ and check (i) if they have an $E_{jj}$ edge and (ii) if they become two new vertices of $G_{l}$ through an neighbor switching event. For the second check, we use their $N_{j}$ triplets. We use Fig. 3a as an example. $N_{j_{1}}=\{g_{1},g_{3},g_{4}\}$ and $N_{j_{2}}=\{g_{2},g_{3},g_{4}\}$. If an edge event happened on $(j_{1},j_{2})$, two new triplets $N_{j^{\prime}_{1}}=\{g_{1},g_{2},g_{3}\}$ and $N_{j^{\prime}_{2}}=\{g_{1},g_{2},g_{4}\}$ must have formed at an intermediate time. We check that if $N_{j^{\prime}_{1}}$ and $N_{j^{\prime}_{2}}$ exist in $G_{l}$, criteria (ii) is satisfied. Thus neighbor-switching events $\mathcal{S}_{E}$ are junction pairs in $\mathcal{U}_{l-1}$ that satisfy (i) and (ii). We remark this method cannot match all the junction vertices. When a junction has been involved in two or more edge or grain elimination events (e.g., Fig. 5a), we fail to detect the current structure. In this case, we mask out (in training) the unmatched junctions and $E_{jj}$ edges. These are typically less than 1% of the total vertices and edges.
Choosing $\Delta z$. Finally, we discuss how to choose $\Delta z$, the distance between $I_{l-1}$ and $I_{l}$. If $\Delta z$ is too small, GrainGNN will need too many steps during inference, and makes the surrogate too expensive. If $\Delta z$ is too large, $I_{l-1}$ and $I_{l}$ will be exceedingly different as the will involve a large number of topological events. With these observations in mind, we choose the number of layers $n_{l}$ such that:
@@
Fig. 4: The LSTM architecture. We use a graph transformer LSTM with a regressor $\mathcal{R}$ and a classifier $\mathcal{C}$—both having an encoder-decoder structure. The classifier and regressor are trained separately. The junction nodes (solid circles) and grain nodes (hollow circles) have vertex features defined in Eq. 2.4 and Eq. 2.5. For each target vertex (red circles), graph transformer operators in Eq. 2.8 aggregates the features of the vertex itself with the neighboring vertices into a hidden vector. The decoder $\mathcal{R}$ then uses Section 2.5 to transform the hidden vector to the target outputs, which are displacements of junctions, area change, and excess volume of grains. For each junction-junction edge, the model $\mathcal{C}$ concatenates the hidden vectors of the two connecting vertices and predicts the probability of the edge event (see Eq. 2.13).

### 31 Jan 2024 v2

@@ 2.1 Problem formulation and high fidelity model
Given the grain microstructure, we compute several quantities of interest (QoIs): number (or percentage) of eliminated grains; grain size distribution; and volume-averaged misorientation. The number of eliminated grains is relevant in epitaxial growth [55]. We compute $n_{G}=n_{G}(z)$, the accumulated grain eliminations from the initial interface location to a height $z$ (see Fig. 2f). Grain size distribution is a common descriptor of grain shape statistics and it affects mechanical properties such as strength and ductility. Grain size $d$ of grain $g$ is defined by its volume-equivalent diameter $d_{g}=(6\mathcal{V}_{g}/\pi)^{1/3}$, where $\mathcal{V}_{g}$ is the grain volume. Fig. 2d shows the probability distributions of the grain size for Fig. 2a and Fig. 2b. Volume-weighted misorientation $\Delta\boldsymbol{\theta}$ quantifies the alignment of the polycrystalline orientation with the prescribed temperature gradient direction $z$. $\Delta\boldsymbol{\theta}=\sum_{g}\mathcal{V}_{g}\boldsymbol{\theta}_{z,g}/\sum_{g}\mathcal{V}_{g}$ and $\boldsymbol{\theta}_{z,g}$ is the angle between the crystal orientation of grain $g$ and the $z$-axis. As shown in Fig. 2f, $\Delta\boldsymbol{\theta}$ decreases as the more aligned grains out-compete the misaligned grains. Due to random initial condition $\boldsymbol{\xi}$, we need to average across several simulations to obtain expectations of the QoIs.
@@ 2.2 GrainGNN
The input to GrainGNN is the initial state $\boldsymbol{\xi}$ of the substrate (see Fig. 2a), process parameters $\boldsymbol{p}=\{G_{z},R_{z}\}$, and a domain size $L=(L_{x},L_{y},L_{z})$. The output of GrainGNN is the grain orientation field $\boldsymbol{\theta}(x,y,z)$ in a domain with size $L$. Notice that here $L_{z}$ is the height of the final polycrystal (as shown in Fig. 2b); for phase field simulations we let $L^{\prime}_{z}>L_{z}$ to avoid the effect of the top boundary. In sections Section 2.3-Section 2.6 we explain the basic ingredients of GrainGNN.
  • Image-to-graph step. We start with a spatial field compression method to reduce the large dimensionality of the phase field data. We define a planar graph $G$ and hand-crafted features $F$ to represent a layer of grain microstructure (see Fig. 3). To define these features from $\{\phi_{\alpha}\}_{\alpha=1}^{n_{g}}$ we first combine the fields to a single image $\boldsymbol{\theta}(x,y,z)$. Then we propose an image-to-graph algorithm $\phi_{\alpha}(x,y,z)\rightarrow(G_{l},F_{l}),\ l=0,1,2,...$ to extract vertices and edges of grains at cross-sections defined at $z_{l}$. The features contain geometric information about the size of the grains and their boundary. The definitions of the image, graph, and features are discussed in Section 2.3.
  • Graph evolution step. We now train a network for the map $(G_{l-1},F_{l-1})\rightarrow(G_{l},F_{l})$. The challenge is that graphs $G_{l-1}$ and $G_{l}$ can be topologically different due to grain and/or edge elimination events. To address this we need to predict both vertex feature evolution $\Delta F$ as well as a set of grain neighbor switching events $\mathcal{S}_{E}$ and a set of grain elimination events $\mathcal{S}_{G}$. We discuss the graph evolution in detail in Section 2.4. The prediction $(G_{l},F_{l})$ is done by first using GrainGNN to predict $\Delta F$ and probabilities of events and then using an algorithm to reconstruct $G_{l}$ and its features.
@@ 2.3 Graph representation of microstructure
Now, let us describe the compressed representation of the microstructure using hand-crafted features. Consider a solution of Eq. 2.1a. From this simulation we sample $n_{l}$ layers of field data. We define a “layer” at height $z_{l}$ as the microstructure at the time $t_{l}$ when the solid-liquid interface (SLI) reaches $z_{l}$ (see Fig. 3a). More precisely, $t_{l}$ is defined as the time at which the lowest point (order in $z$ coordinate) at the SLI reaches $z_{l}$.
@@
We solve Eq. 2.1a to generate training data and verify the predictions of the surrogate. Although a simulation can have thousands of time steps, we subsample to $n_{l}$ planes using equispaced sampling for $z_{l}$ with height increment $\Delta z$, so $z_{l}=l\Delta z$. Typically we use $n_{l}=20$. We discuss the choice of $\Delta z$ and $n_{l}$ in Section 2.4.
@@
Once $I_{l}$ have been identified, we extract graph $(G_{l},F_{l}$ (Fig. 3c) as follows. $G(V,E)$ is an undirected graph where $V$ and $E$ denote the sets of vertices and edges respectively. In GrainGNN, $G$ has two types of vertices $V=V_{g}\cup V_{j}$. $V_{g}=\{{g}_{1},g_{2},...,{g_{n_{g}}}\}$ represents grains. $V_{\boldsymbol{j}}=\{{j}_{1},{j}_{2},...,{j_{n_{j}}}\}$ are junction vertices, where $n_{j}$ is the number of junctions. A junction vertex refers to a point in image $I_{l}$ that has three grain neighbors (e.g., $j_{1}$ in Fig. 3d). Notice that the junction vertices are just the vertices of the Voronoi diagram. $G$ has two types of undirected edges $E=E_{jj}\cup E_{jg}$. $E_{jj}$ are the edges between junction vertices and can be thought of as a representation of actual inter-grain boundaries on the $x$-$y$ plane. $E_{jg}$ are the junction-grain edges (dashed lines in Fig. 3e).
@@
As we will see later we need to define a maximum degree (number of edges) for a junction vertex. The maximum number of edges for a junction vertex depends on the physical setting of the problem. In rapid solidification of metals, the grain boundary anisotropy is relatively weak compared to kinetic anisotropy [56], and a triple junction is more stable than a quadruple junction [19]. Thus we only consider triple junctions in this work, which means a junction vertex can only connect to the other three junction vertices and three grain vertices. In this case, $n_{j}=2n_{g}$, $|E_{jj}|=3n_{g}$, and $|E_{jg}|=6n_{g}$. Our handling of the graph topological changes discussed in Section 2.4 is based on this simplification of the graph structure.
@@
Let’s now discuss how to identify vertices and edges from $I_{l}$. First, we identify the junction locations $(x_{j}$, $y_{j})$. For each pixel $(x,y)$ of $I_{l}$, we inspect its eight neighbors each associated with a grain index. If there are three distinctive grain indices in eight neighbors, we consider pixel $(x,y)$ to be a junction pixel, for example, $j_{1}$ in Fig. 3d. Every junction $j$ is associated with a unique triplet of grain indices $N_{j}=\{g_{1},g_{2},g_{3}\}$. Commonly multiple adjacent pixels will have the same triplet and we only keep one of them [^2] If two or more pixels have the same triplet $\{g_{1},g_{2},g_{3}\}$. We count the occurrences of each grain index in the eight neighbors of each pixel and use the one pixel whose index occurrences are more even. For example, for pixel 1 the occurrences are $\{3,3,2\}$ and for pixel 2 the occurrences are $\{4,3,1\}$ and we choose pixel 1.. Then using the triplet indices $N_{j}$ we add three edges to the $E_{jg}$ set: e.g., for $j_{1}$, the $E_{jg}$ edges are $({j_{1}},g_{1})$, $({j_{1}},g_{2})$, $({j_{1}},g_{3})$. We create $E_{jj}$ edges by $E_{jj}=\{(j_{1},j_{2}):N_{j_{1}}\cap N_{j_{2}}=2,\forall j_{1},j_{2}\in G\}$. For example in Fig. 3e, edge $(j_{1},j_{2})$ exists because $j_{1}$ and $j_{2}$ share the same neighboring grains $g_{1}$ and $g_{2}$.
@@
We clarify that these are input features to the LSTM and serve as a reduced representation of the grain structure. They are not the network hidden features, which are of course determined during training. We define vertex and edge features as follows:
@@
    where $x_{j}$, $y_{j}$, and $z_{l}$ are the 3D coordinates of a junction vertex; $\Delta x_{j}$ and $\Delta y_{j}$ are in-plane displacements of junctions with respect to their locations in the previous layer $G_{l-1}$, i.e., $\Delta x_{j}=x_{j,l}-x_{j,l-1},\Delta y_{j}=y_{j,l}-y_{j,l-1}$. $\Delta z=z_{l}-z_{l-1}$ is the distance between two sampled images. For $l=0$, we set $\Delta x_{j}=\Delta y_{j}=\Delta z=0$. $G_{z}$ and $R_{z}$ here are constants; they’re repeated at each vertex and we allow them to vary for each junction location.
@@
    where $x_{g}$, $y_{g}$, and $z_{l}$ are the coordinates of the grain vertices; $s_{g}$ is the grain cross-sectional area; and $v_{g}$ is the grain excess volume above the interface height $z_{l}$ [17]. For $x_{g}$, we average the coordinates of its junction neighbors $x_{g}=\sum_{k\in N_{g}}x_{j,k}/|N_{g}|$. $N_{g}$ is the set of junctions connected to grain $g$. $\boldsymbol{\theta}_{z}$ is the angle between the $z$-axis and the preferred growth direction ¡100¿; $\boldsymbol{\theta}_{x}$ is the angle between the $x$-axis and the projection of ¡100¿ direction on the $x$-$y$ plane. $\Delta s_{g}=s_{g,l}-s_{g,l-1}$ is the change of cross-sectional area; $\Delta s_{g}=0$ at $l=0$. In Eq. 2.5 and Eq. 2.4, $\boldsymbol{f}^{0}_{j}$ and $\boldsymbol{f}^{0}_{g}$ are constant vectors to normalize features to [0, 1]. $\boldsymbol{f}^{0}_{j}=\begin{bmatrix}L_{x},L_{y},L_{z},G_{\mathrm{max}},R_{\mathrm{max}},L_{x},L_{y},L_{z}\end{bmatrix}$ and $\boldsymbol{f}^{0}_{g}=\begin{bmatrix}L_{x},L_{y},L_{z},L_{x}L_{y},L_{x}L_{y}L_{z},1,1,1,1,L_{x}L_{y},L_{z}\end{bmatrix}$. $G_{\mathrm{max}}$ and $R_{\mathrm{max}}$ are maximum values of $G_{z}$ and $R_{z}$ in the training data.
@@
Fig. 4: Topological events with the update of graph connectivity. (a) A neighbor-switching event requires changes of the graph topology: an $O_{E}$ operation. Grains $g_{3}$ and $g_{4}$ initially are neighbors. After applying $O_{E}$, $g_{1}$ and $g_{2}$ are neighbors. From \@slowromancapi@ to \@slowromancapii@, $j_{1},j_{2}$ merge to a quadruple junction; from \@slowromancapii@ to \@slowromancapiii@, the quadruple junction splits into two triple junctions. This event replaces edges $(j_{1},j_{5})$, $(j_{2},j_{4})$ (in red color) with the edges $(j^{\prime}_{2},j_{5})$, $(j^{\prime}_{1},j_{4})$ (in blue color). The total number of edges remains the same. (b) A grain elimination event. From \@slowromancapi@ to \@slowromancapii@, a neighbor-switching event occurs on $(j_{1},j_{2})$; $g_{3}$ becomes a 2-side grain. \@slowromancapii@ to \@slowromancapiii@ is an $O_{G}$ operation. It removes $g_{3}$, $j^{\prime}_{1}$, $j_{3}$, a couple of edges, and replace them with $(j^{\prime}_{2},j_{4})$.
@@ 2.4 Graph evolution
Grain neighbor switching events: Grain neighbor switching happens when a grain edge becomes smaller and eventually disappears. A disappearing edge creates a four-edge junction (Fig. 4a). Since in our PDE model four-edge junctions are unstable, a new edge starts growing and the four-junction becomes two triple junctions. This new edge causes the neighbor to switch. Physically, a neighbor-switching event involves four grains, where two grains lose one face and the other two grains gain one face. From \@slowromancapi@ to \@slowromancapii@, grain $g_{1}$ and $g_{2}$ push the vertices $j_{1}$ and $j_{2}$ to move toward each other. At some intermediate time, $j_{1}$ and $j_{2}$ merge to form a quadruple junction. Then from \@slowromancapii@ to \@slowromancapiii@, the quadruple junction separates into two new junctions $j^{\prime}_{1}$ and $j^{\prime}_{2}$, which replace $j_{1}$ and $j_{2}$. This event changes several edges, grain $g_{1}$ and $g_{2}$ grain an additional junction-grain edge; $g_{3}$ and $g_{4}$ lose one junction-grain edge; $(j^{\prime}_{2},j_{5})$ and $(j^{\prime}_{1},j_{4})$ replace $(j_{1},j_{5})$ and $(j_{2},j_{4})$; the total number of edges remains the same. The grain neighbor-switching event is characterized by the removal of $(j_{1},j_{2})$. We will refer to a grain neighbor switching event as an “edge” event in the following discussion.
@@
Matching $G_{l-1}$ and $G_{l}$ and handling elimination events: Our graph update requires computing $\Delta F,\mathcal{S}_{E},\mathcal{S}_{G}$. We define
@@
Neighbor-switching events are harder to handle: they result in junctions whose triplets $N_{j}$ are different in $G_{l}$ and $G_{l-1}$. Let $\mathcal{U}_{l-1}$ be the junction vertices in $G_{l-1}$ but not matched in $G_{l}$, $\mathcal{U}_{l-1}=\{j:j\in G_{l-1}\wedge j\notin G_{l}\}$. We enumerate every two junctions in $\mathcal{U}_{l-1}$ and check (i) if they have an $E_{jj}$ edge and (ii) if they become two new vertices of $G_{l}$ through a neighbor switching event. For the second check, we use their $N_{j}$ triplets. We use Fig. 4a as an example. $N_{j_{1}}=\{g_{1},g_{3},g_{4}\}$ and $N_{j_{2}}=\{g_{2},g_{3},g_{4}\}$. If an edge event happened on $(j_{1},j_{2})$, two new triplets $N_{j^{\prime}_{1}}=\{g_{1},g_{2},g_{3}\}$ and $N_{j^{\prime}_{2}}=\{g_{1},g_{2},g_{4}\}$ must have formed at an intermediate time. We check that if $N_{j^{\prime}_{1}}$ and $N_{j^{\prime}_{2}}$ exist in $G_{l}$, criteria (ii) is satisfied. Thus neighbor-switching events $\mathcal{S}_{E}$ are junction pairs in $\mathcal{U}_{l-1}$ that satisfy (i) and (ii). We remark this method cannot match all the junction vertices. When a junction has been involved in two or more edge or grain elimination events (e.g., Fig. 6a), we fail to detect the events. In this case, we mask out (in training) the unmatched junctions and $E_{jj}$ edges. These are typically less than 1% of the total vertices and edges.
Choosing $\Delta z$: Finally, we discuss how to choose $\Delta z$, the distance between $I_{l-1}$ and $I_{l}$. If $\Delta z$ is too small, GrainGNN will need too many steps during inference, and make the surrogate too expensive. If $\Delta z$ is too large, $I_{l-1}$ and $I_{l}$ will be exceedingly different as they will involve a large number of topological events. With these observations in mind, we choose the number of layers $n_{l}$ such that:
@@
Fig. 5: The LSTM architecture. We use a graph transformer LSTM with a regressor $\mathcal{R}$ and a classifier $\mathcal{C}$—both having an encoder-decoder structure. The classifier and regressor are trained separately. The junction nodes (solid circles) and grain nodes (hollow circles) have vertex features defined in Eq. 2.4 and Eq. 2.5. For each target vertex (red circles), graph transformer operators in Eq. 2.8 aggregate the features of the vertex itself with the neighboring vertices into a hidden vector. The decoder $\mathcal{R}$ then uses Section 2.5 to transform the hidden vector to the target outputs, which are displacements of junctions, area change, and excess volume of grains. For each junction-junction edge, the model $\mathcal{C}$ concatenates the hidden vectors of the two connecting vertices and predicts the probability of the edge event (see Eq. 2.13).